OpenStudy (anonymous):

What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10−6 cm/s B. 8.889 × 10−5 cm/s C. 1.152 × 10−5 cm/s D. 1.152 × 10−7 cm/s

OpenStudy (anonymous):

@Allofus

OpenStudy (anonymous):

Do you mean that m/h is meters/hour and thx

OpenStudy (anonymous):

I think it means miles per hour but im not sure

OpenStudy (anonymous):

It might be meters

OpenStudy (anonymous):

if it were miles it would be mph but we can check both since its multiple choice

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hmmm... does 8.889 × 10−6 cm/s mean 8.889 × 4 cm/s or \[8.889 × 10^{-6}\]

OpenStudy (anonymous):

Oh sorry, im new to Open Study, all the x10^(number here)

OpenStudy (anonymous):

so yes it is 10^-6

OpenStudy (anonymous):

@Allofus you still there?

OpenStudy (anonymous):

Yep so sorry I gotta eat real quick

OpenStudy (anonymous):

...

OpenStudy (anonymous):

OK 3.2*10^-4 * 60sec/1min = 0.0192 0.0192 *100 = 1.92 Doesn't work

OpenStudy (anonymous):

Hmmm...

OpenStudy (anonymous):

ok.. i found that 3.2 × 10−4 = 0.08888888888889

OpenStudy (anonymous):

*10^-4

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

I think it is either A or B

OpenStudy (anonymous):

@Allofus do you know?

OpenStudy (anonymous):

WAIT can you do a drawing what the equation looks like

OpenStudy (anonymous):



OpenStudy (anonymous):

Then its equal to 0.00032

OpenStudy (anonymous):

So is it A B C Or D? I'm confused

OpenStudy (anonymous):

Me too

OpenStudy (anonymous):

@phi help help help

OpenStudy (anonymous):

I've used mathway for this problem, not much help :/

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

I just chose B. last question: On average, a commercial airplane burns about 4.0 × 103 mL of gasoline per second. There can be as many as 5 × 103 airplanes in the air in the United States at any given time. About how many gallons of gasoline are burned per hour by commercial airplanes in flight in the United States? (1 gal = 3,785.41 mL). 1.9 × 10 gal

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

jut wondering; what grade is this for??

OpenStudy (anonymous):

8th grade

OpenStudy (anonymous):

@thundernova308 did u check the link

OpenStudy (amistre64):

3.2 × 10−4 m/h change m and h to cm and s appropriately

OpenStudy (amistre64):

m = 100 cm h = 3600 s

OpenStudy (anonymous):

@sheelo_mughal , yes thanks for that btw, but sadly not much help

OpenStudy (amistre64):

3.2 (10^-4) ---------- cm/s 36

OpenStudy (anonymous):

@amistre64 is this for the last question I post?

OpenStudy (amistre64):

that should be self evident .....

OpenStudy (amistre64):

i was responding to your orignal post, which is what i was tagged in for i believe

OpenStudy (anonymous):

I'm guessing no :/ I answered it already, can you help with the last question

OpenStudy (amistre64):

new questions should be asked as new questions. please close out old answered questions.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Wait, we only guessed the original question!!

OpenStudy (anonymous):

but why \[3.2\times 10^-4=0.00032\] am i right @Allofus

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so \[3.2\times 10^-4m/h=8.9\times 10^-17cm/s\]